OpenStudy (anonymous):

what is the equation of a circle whose center is at the origin and whose radius is 9

OpenStudy (gorv):

(x-a)^2+(y-b)^2=r^2 (a,b) is centre and r= radius

OpenStudy (gorv):

at origin a=0 and b=0 and radius is given plug in all the values

OpenStudy (anonymous):

sooo.... it would be x^2+y^2=3

OpenStudy (gorv):

lol r=9 9^2= x^2+y^2=9^2

OpenStudy (anonymous):

so its X^2+Y^2=9 lol

OpenStudy (gorv):

9^2=81 not 9